OpenStudy (anonymous):

Completely factor the polynomial: 2x2 − 30x + 88

OpenStudy (anonymous):

2x^2 -30x +88 take 2 out as it is the common factor of the 3 terms 2(x^2-15x+44) (-4)+(-11)=-15 and (-4)(-11)=+44 therefore, 2(x^2-15x+44) =2(x-4)(x-11)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

you are welcome